# Hybrid Two-Stage Reconstruction of Multiscale Subsurface Flow with Physics-informed Residual Connected Neural Operator

### v1

@@ Abstract
keywords:
Multiscale Modeling; Subsurface Fluid Flow Simulation; Mixed Generalized Multiscale Finite Element Method; PDE Solver; Physics-informed Neural Operator; Two-stage Method
journal: Journal of Computational Physics
[…]
[SMP]organization=School of Mathematics and Physics, Xi’an Jiaotong-Liverpool University, postcode=215123, city=Suzhou, country=China
@@ 1 Introduction
After decades of development, traditional numerical solvers (finite difference methods, finite element methods, etc.) have demonstrated outstanding effectiveness and robustness in solving complex partial differential equations (PDEs)[1]. However, when dealing with problems that require high iterations or parameter inversion, these traditional methods do not have a clear advantage in terms of efficiency[2]. Taking the problem of fluid flow in porous media, especially in fractured porous media with high contrast and heterogeneity, as an example, the solution process requires consideration of complex fracture systems[3]. The fractures in the system vary significantly in size and direction, causing the problem to span multiple scales. In such multiscale systems, using large-scale models may reduce the accuracy of the model due to the neglect of small fractures, while small-scale computations significantly increase computational costs. Therefore, the introduction of multiscale techniques is essential, providing an effective solution for simulating fluid flow in heterogeneous porous media[4].
Multiscale techniques have widespread applications in solving fluid flow problems. Common multiscale methods include the Multiscale Finite Volume Method (MsFVM)[5, 6], Multiscale Finite Element Method (MsFEM)[7, 8], and Mortar Multiscale Method[9, 10]. These methods solve problems by applying precomputed multiscale basis functions on coarse grids. Jenny et al.[5]. proposed a finite volume method based on a flux-continuous finite difference scheme, which simplifies the flow problem by solving small-scale problems locally and constructing finite large-scale transfer coefficients. The method also introduces new basis function designs to ensure mass conservation when reconstructing the fine-scale velocity field. This method can reconstruct a velocity field that satisfies local fine-scale mass conservation, making the global solution at the fine scale more accurate. For the multiscale finite element method, it introduces multiple basis functions to enhance the model’s ability to handle complex problems. At the same time, spectral reduction techniques are employed to maintain model accuracy while reducing computational complexity. Building on this, the mixed Generalized Multiscale Finite Element Method[11] combines multiscale basis functions with direct velocity field solutions to ensure local conservation, particularly suitable for simulating high-contrast and multiphase flow. By selecting an appropriate offline basis function space, this method reduces the computational burden during online calculations while maintaining high accuracy using fewer basis functions.
@@
With the development of artificial intelligence (AI) technology, neural networks have demonstrated remarkable capabilities learning complex mappings, and they exhibit computational speeds significantly faster than traditional numerical methods. Therefore, training neural network operators[12] to map random parameters to the solutions of PDEs has become an indispensable method in fields like scientific computing. The most common neural operators are finite operators, such as Convolutional Neural Networks (CNNs)[13], which extract features via local convolutional kernels. For instance, Choubineh et al.[14] used a convolutional neural network to build a model for reconstructing multiscale basis functions and pressure field solutions in mixed GMsFEM. Through feature learning with convolutional operators, the model can map random parameters to the equation’s solution through complex linear and nonlinear transformations. Additionally, their deep ensemble model[15] integrates shortcut connections and linearly fits the results of different network paths to approximate the corresponding solution. This model achieves extremely high accuracy in solving the pressure field of Darcy’s equation. In contrast to finite operators, infinite operators do not discretize the input parameters and output solutions. A typical example of an infinite operator is the Fourier Neural Operator (FNO)[16, 17], which uses Fourier transform to map the input function parameters from the spatial domain to the frequency domain, decomposing the information into multiple Fourier modes with different frequencies, and learning the weights of these modes in the frequency domain.
@@
For the training of neural networks, there are two common methods: data-driven method and Physics-informed method. In the data-driven approach, based on the data requirements of deep learning technology, traditional PDE solvers are usually used to simulate the solution functions of a large number of different parameter functions, and then these data are used for the training of neural networks. While this training method can achieve desired results, it usually requires a large amount of data to train and can cause the trained operators to violate the laws of physics. In contrast, physics-informed methods[24, 25] ensure the superiority of local physical constraints and require fewer samples for training. Raissi et al.[26] proposed the Physics-informed Neural Networks (PINNs), which embedded the physical information of PDE into the training process of the neural network through automatic differentiation technology. This technique is achieved by incorporating physical losses into the training objectives of the optimizer.
In this paper, a two-stage neural network operator is proposed to solve the Darcy flow problem step by step in the mixed GMsFEM framework: firstly, the multiscale basis function defined on the coarse grid is reconstructed through data-driven model training, and then the pressure field is solved through physics-informed model training. Fig.1 shows the flow chart of our study. On the left is the process of building the entire model, which applies to two stages. The right side shows the general architecture of the two-stage model training.
The remaining sections of this paper will address different aspects of the study. Sec.2 will mainly explain the dataset we used; Sec.3 and 4 will describe the work carried out in the two stages, including the methods used and the reconstruction tasks that need to be implemented. Sec.5 will explain the results and evaluation of the two-stage models. Sec.6 will summarize the main findings of this study and discuss its practical implications and future research directions. The appendix will provide additional details on the theoretical framework of the mixed GMsFEM method applied to our problem and present examples of our results. In the A, we will introduce the knowledge of mixed GMsFEM for Darcy’s equation. And the BLABEL:appendix:examples will illustrate some more examples of the results of our proposed method.
@@ 2 Dataset Description
[…]
@@
Fig.2 shows an example of the dataset in our research, from left to right: high-contrast permeability field, reconstructed multiscale basis functions contour plot (reshaped to $30\times 30$), and a pressure field. In a multiscale system, the fine mesh is configured as a $30\times 30$ homogenized grid, and the coarse grid is a $10\times 10$ grid, that is, each coarse mesh contains $3\times 3$ fine grid. In the grid system we define, the coarse grid may contain all or part of the fractures, which can be parallel or intersecting. In the parameter setting of permeability, the matrix permeability $K_{m}$ is between { 1, 2, 3, 4, 5 } millidarcy values, the fracture permeability $K_{f}$ is between { 500, 750, 1000, 1250, 1500, 1750, 2000 } millidarcy values, and the fracture number is between 1 and 25 random integers. The combination of three random parameters can generate 875 separate combinations for all for (permeability field, multiscale basis function) pairs. By iteratively generating these combinations, a total of 177,800 samples were generated. Considering the possibility of generating duplicate data, we removed 6,537 duplicates after removing the duplications. The dataset was divided into a 6:2:2 ratio of 102,757 training samples, 34,252 validation samples, and 34,254 test samples. The pressure field is a two-dimensional matrix obtained directly by mixed GMsFEM. For stage2, there are totally 1,700 samples envolved in physics-informed model training.
@@ 3 Multiscale Preconditioner for Reconstruction of Basis Functions : Stage 1
3.1 Domain Transformation-based Preconditioner
The first step in this phase of the model involves using a Fourier Neural Operator-based neural network for data preprocessing[16], aimed at extracting frequency information inherent in the data. This preconditioner employs the Fourier integral operator $\mathcal{K}$ to transform the input $x$ into a spectral representation, a process that captures both local and global dependencies, thereby gaining insights into the spatial relationships within multiscale basis functions.
@@ 3.1 Domain Transformation-based Preconditioner
here, $\psi(x,\xi)=\exp(2\pi i\langle x,\xi\rangle\in L(D)$ is the Fourier basis function, $\xi$ represents the frequency mode used for nonlinear transformations within the network, indicating the Fourier modes to be retained. $\mathcal{T}$ is a uniform grid sampled from the distribution $\mu$.
@@
$\hat{x}(\tilde{\xi})=\int_{\Omega}x\exp\left(-2\pi i\langle x,\tilde{\xi}\rangle\right)$ (3)
@@
$\hat{y}(\tilde{\xi})=\tilde{\mathcal{K}}(\tilde{\xi})\cdot\hat{x}(\tilde{\xi})$ (4)
where $\tilde{\mathcal{K}}(\cdot)$ is the learnable kernel in the frequency domain, allowing the interactions under different spatial scales.
@@
[…]
@@ 3.2 Parallel Hierarchical Network for Multiscale Information
In this section, we will introduce the second part of Stage 1. After the preconditioner, data will be fed into a convolutional neural network (CNN) with a parallel architecture for further feature learning.
Although multiscale basis functions are defined on a coarser grid system, valuable information is still present in the finer grid system. To capture this information, we have designed convolutional filters of two different scales for information learning, which do not alter the size of the data. Our multiscale network is implemented through two parallel convolutional modules. In this design, the output of the preprocessor will pass through a larger-scale network ($3\times 3$) and a smaller-scale network ($1\times 1$) simultaneously. The $1\times 1$ convolutional filter does not involve complex data operations and can be considered an expansion of the original data channels.
@@ 3.3 Evaluation Metircs for this stage
Although we propose a two-stage approach, it is imperative to validate the experimental outcomes of the model from the first stage before proceeding to the second. At this juncture, our focus is particularly on the accuracy of the model’s fit to the data. Consequently, we opt to evaluate the model using the Mean Squared Error (MSE) and the Coefficient of Determination (R2).
@@
R2 represents the proportion of variance in the dependent variable that is explained by the independent variables, serving as an indicator of the model’s fit quality.
[…]
@@ 4 Physical Connected Operator Solving Pressure Fields : Stage 2
This section will detail the second phase of our methodology: a PINN-Transformer architecture capable of reconstructing pressure fields from small sample data, featuring residual connections. In this architecture, the permeability field $k$ and the multiscale basis functions $bf$ obtained from the first stage will serve as input data to learn the pressure solution $p(k,bf)$ of Darcy’s equation. In other words, we aim to develop a parameterized deep learning model $\mathcal{G}$ that facilitates the mapping from input to output:
$\mathcal{G}:\ \mathbb{R}^{2\times 2}\mapsto\mathbb{R}^{2},\ \ \ (k,bf)\mapsto p_{\theta}(k,bf)$ (10)
@@ 4.1 Physics-informed Operator with Non-neighbored Accelerator
PGDL methods combine the powerful nonlinear mapping capabilities of deep learning with prior knowledge of physical laws, thereby significantly enhancing the accuracy, stability, and interpretability of predictive models. The Physics-Informed Neural Network (PINN) serves as a specific implementation of PGDL, embedding physical information into the model through residual learning.
@@
$\mathcal{R}(x;\theta)=-\nabla\cdot\left(k(x)\nabla p_{\theta}(x)\right)-f(x)=-\left(\nabla k(x)\cdot\nabla p_{\theta}(x)+k(x)\Delta p_{\theta}(x)\right)-f(x)$ (11)
@@
To embed physical residuals into a neural network, we utilize a loss function to achieve this objective. Traditional neural networks typically employ the Mean Squared Error (MSE) loss for regression tasks or the Binary Cross-Entropy (BCE) loss for binary classification tasks. However, in our research, we need to incorporate physical information into the loss function to ensure that the model’s predictions adhere to the laws of physics. Additionally, we must consider the loss associated with the data itself to enhance the model’s fit to the actual observed data.
@@
$\mathcal{L}_{\text{data}}(\theta)=\frac{1}{N}\sum_{j=1}^{N}\left[p_{\theta}(x_{j})-p_{\text{true}}(x_{j})\right]^{2}$ (12)
@@
$\mathcal{L}_{\text{physics}}=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left[\mathcal{R}(x_{i};\theta)\right]^{2}$ (14)
@@
Fig.4-stage2-local information extractor illustrates the operator architecture we propose, which consists of symmetrically distributed V-shaped convolutional layers and embeds physical information through residual learning during backpropagation. The increase in model complexity may lead to a higher risk of overfitting, which is undesirable. To address this, in addition to incorporating Dropout layers to mitigate the potential for overfitting, we have also introduced L2 regularization (ridge regression) into the model, which adds a regularization term to the loss function:
@@
As the network architecture becomes increasingly complex, a significant loss of data features may occur during the transmission of information between layers, potentially leading to the gradient vanishing problem. Additionally, a large number of parameters can substantially prolong the training time of the model. To effectively expedite the model training process, we have incorporated accelerators based on residual connections into the network architecture, aiming to mitigate these issues and enhance training efficiency.
@@
where $\mathcal{G}^{(b)}$ represents the original mapping function from layer $(b-1)$ to layer $b$, a process that is independent of the skip connection.
These non-adjacent skip connections act as ”shortcuts” from earlier layers to later layers in the computational graph, creating a multipath structure. During forward propagation, features from layer $a$ can directly influence the output of layer $b$ without undergoing the nonlinear transformations of the intermediate layers. This design ensures that the feature information extracted by the early layers is not easily diminished or lost during deep propagation.
Assuming the existence of a skip connection in the network that directly jumps from the input layer to the L-th layer:

### v2

@@ Abstract
Keywords: Multiscale Modeling; Subsurface Fluid Flow Simulation; Mixed Generalized Multiscale Finite Element Method; PDE Solver; Physics-informed Neural Operator; Two-stage Method
@@ 1 Introduction
After decades of development, traditional numerical solvers (finite difference methods (FDM), finite element methods (FEM), etc.) have demonstrated outstanding effectiveness and robustness in solving complex partial differential equations (PDEs)[1]. However, when dealing with problems that require high iterations or parameter inversion, these traditional methods do not have a clear advantage in terms of efficiency[2]. Taking the problem of fluid flow in porous media, especially in fractured porous media with high contrast and heterogeneity, as an example, the solution process requires consideration of complex fracture systems[3]. The fractures in the system vary significantly in size and direction, causing the problem to span multiple scales. In such multiscale systems, using large-scale models may reduce the accuracy of the model due to the neglect of small fractures, while small-scale computations significantly increase computational costs. Therefore, the introduction of multiscale techniques is essential, providing an effective solution for simulating fluid flow in heterogeneous porous media[4].
Multiscale techniques have widespread applications in solving fluid flow problems. Common multiscale methods include the Multiscale Finite Volume Method (MsFVM)[5, 6], Multiscale Finite Element Method (MsFEM)[7, 8], and Mortar Multiscale Method[9, 10]. These methods solve problems by applying precomputed multiscale basis functions on coarse grids. Jenny et al.[5]. proposed a finite volume method based on a flux-continuous finite difference scheme, which simplifies the flow problem by solving small-scale problems locally and constructing finite large-scale transfer coefficients. The method also introduces new basis function designs to ensure mass conservation when reconstructing the fine-scale velocity field. This method can reconstruct a velocity field that satisfies local fine-scale mass conservation, making the global solution at the fine scale more accurate. For the multiscale finite element method, it introduces multiple basis functions to enhance the model’s ability to handle complex problems. At the same time, spectral reduction techniques are employed to maintain model accuracy while reducing computational complexity. Building on this, the mixed Generalized Multiscale Finite Element Method[11] (mixed GMsFEM) combines multiscale basis functions with direct velocity field solutions to ensure local conservation, particularly suitable for simulating high-contrast and multiphase flow. By selecting an appropriate offline basis function space, this method reduces the computational burden during online calculations while maintaining high accuracy using fewer basis functions.
@@
With the development of artificial intelligence (AI) technology, neural networks have demonstrated remarkable capabilities in learning complex mappings, and they exhibit computational speeds significantly faster than traditional numerical methods. Therefore, training neural network operators[12] to map random parameters to the solutions of PDEs has become an indispensable method in fields like scientific computing. The most common neural operators are finite operators, such as Convolutional Neural Networks (CNNs)[13], which extract features via local convolutional kernels. For instance, Choubineh et al.[14] used a convolutional neural network to build a model for reconstructing multiscale basis functions and pressure field solutions in mixed GMsFEM. Through feature learning with convolutional operators, the model can map random parameters to the equation’s solution through complex linear and nonlinear transformations. Additionally, their deep ensemble model[15] integrates shortcut connections and linearly fits the results of different network paths to approximate the corresponding solution. This model achieves extremely high accuracy in solving the pressure field of Darcy equation. In contrast to finite operators, infinite operators do not discretize the input parameters and output solutions. A typical example of an infinite operator is the Fourier Neural Operator (FNO)[16, 17], which uses Fourier transform to map the input function parameters from the spatial domain to the frequency domain, decomposing the information into multiple Fourier modes with different frequencies, and learning the weights of these modes in the frequency domain.
@@
For the training of neural networks, there are two common methods: data-driven method and physics-informed method. In the data-driven approach, based on the data requirements of deep learning technology, traditional PDE solvers are usually used to simulate the solution functions of a large number of different parameter functions or fields, and then these data are used for the training of neural networks. While this training method can achieve desired results, it usually requires a large amount of data to train and can cause the trained operators to violate the laws of physics. In contrast, physics-informed methods[24, 25] ensure the superiority of local physical constraints and require fewer samples for training. Raissi et al.[26] proposed the Physics-informed Neural Networks (PINNs), which embedded the physical information of PDE into the training process of the neural network through automatic differentiation technology. This technique is achieved by incorporating physical losses into the training objectives of the optimizer.
In this paper, a two-stage neural network operator is proposed to solve the Darcy flow problem step by step in the mixed GMsFEM framework: firstly, the multiscale basis function defined on the coarse grid is computed through data-driven model training, and then the pressure field is solved through physics-informed model training. Fig.1 shows the flow chart of our study. On the left is the process of building the entire model, which applies to two stages. The right side shows the general architecture of the two-stage model training.
The remainder of this paper will address different aspects of the study. Sec.2 will mainly explain the dataset we used; Sec.3 and 4 will describe the work carried out in the two stages, including the methods used and the reconstruction tasks that need to be implemented. Sec.5 will explain the results and evaluation of the two-stage models. Sec.6 will summarize the main findings of this study and discuss its practical implications and future research directions. The appendix will provide additional details on the theoretical framework of the mixed GMsFEM method applied to our problem and present examples of our results. In the A, we will introduce the knowledge of mixed GMsFEM for Darcy’s equation. And the BLABEL:appendix:examples will illustrate some more examples of the results of our proposed method.
@@ 2 Dataset Description
In the two-stage training process, we used data-driven and physics-informed methods, respectively. This means that in the first stage, a large amount of data is required for model training. In this stage, we parameterize random fields with specific covariance properties (high-contrast permeability fields) using Karhunen-Loève Expansion[27]. The computational domain is defined as $\Omega=[0,1]^{2}$, and the corresponding multiscale basis functions have values ranging from -1 to 1 ($1<\text{bf values}<1$). The first basis function (basis1) is a piecewise constant, and we do not need to reconstruct this data using deep learning (see Fig.3). Therefore, the multiscale basis functions involved in the training are basis2, basis3, basis4, and basis5.
@@
Fig.2 shows an example of the dataset in our research, from left to right: high-contrast permeability field, reconstructed multiscale basis functions contour plot (reshaped to $30\times 30$), and a pressure field. In a multiscale system, the fine grid is configured as a $30\times 30$ homogenized grid, and the coarse grid is a $10\times 10$ grid, that is, each coarse mesh contains $3\times 3$ fine grid. In the grid system, the coarse grid may contain all or part of the fractures, which can be parallel or intersecting. In the parameter setting of permeability, the matrix permeability $K_{m}$ is between { 1, 2, 3, 4, 5 } millidarcy values, the fracture permeability $K_{f}$ is between { 500, 750, 1000, 1250, 1500, 1750, 2000 } millidarcy values, and the fracture number is between 1 and 25 random integers. The combination of three random parameters can generate 875 separate combinations for all for (permeability field, multiscale basis function) pairs. By iteratively generating these combinations, a total of 177,800 samples were generated. Considering the possibility of generating duplicate data, we removed 6,537 duplicates after removing the duplications. The dataset was divided into a 6:2:2 ratio of 102,757 training samples, 34,252 validation samples, and 34,254 test samples. The pressure field is a two-dimensional matrix obtained directly by mixed GMsFEM. For stage2, there are totally 1,700 samples envolved in physics-informed model training.
@@ 3 Multiscale Preconditioner for Reconstruction of Basis Functions: Stage 1
3.1 Domain Transform-based Feature Extraction
The first step in this phase of the model involves using a Fourier Neural Operator-based neural network for the first stage feature extraction[16], aimed at extracting frequency information inherent in the data and filtering redundant information. This part employs the Fourier integral operator $\mathcal{K}$ to transform the input $x$ into a spectral representation, a process that captures both local and global dependencies, thereby gaining insights into the spatial relationships within multiscale basis functions.
@@ 3.1 Domain Transform-based Feature Extraction
here, $\psi(x,\xi)=\exp(2\pi i\langle x,\xi\rangle)\in L(D)$ is the Fourier basis function, $\xi$ represents the frequency mode used for nonlinear transformations within the network, indicating the Fourier modes to be retained. $\mathcal{T}$ is a uniform grid sampled from the distribution $\mu$.
@@
[…]
@@
[…]
@@
$y=\int_{\tilde{\xi}}\hat{y}\exp\left(2\pi i\langle x,\tilde{\xi}\rangle\right)d\tilde{\xi}$ (5)
@@ 3.2 Parallel Hierarchical Network for Multiscale Information
In this section, we will introduce the second part of Stage 1. After the frequency-based learning, data will be fed into a CNN framework with a parallel architecture for further feature learning.
Although multiscale basis functions are defined on a coarser grid system, valuable information is still present in the finer grid system. To capture this information, we have designed convolutional filters of two different scales for information learning, which do not alter the size of the data. Thanks to the idea of oversampling, our multiscale network is implemented through two parallel convolutional modules. In this design, the output of the preprocessor will pass through a larger-scale network ($3\times 3$) and a smaller-scale network ($1\times 1$) simultaneously. The $1\times 1$ convolutional filter does not involve complex data operations and can be considered an expansion of the original data channels.
@@ 3.3 Evaluation Metircs for this stage
Although we propose a two-stage approach, it is imperative to validate the experimental outcomes of the model from the first stage before proceeding to the second. At this juncture, our focus is particularly on the accuracy of the model’s fit to the data. Consequently, we opt to evaluate the model using the Mean Squared Error (MSE) and the Coefficient of Determination ($\text{R}^{2}$).
@@
$\text{R}^{2}$ represents the proportion of variance in the dependent variable that is explained by the independent variables, serving as an indicator of the model’s fit quality.
$\text{R}^{2}=1-\frac{\sum_{i=1}^{N}\left(y_{i}-\hat{y}_{i}\right)^{2}}{\sum_{i=1}^{N}\left(y_{i}-\bar{y}\right)^{2}}$ (9)
@@ 4 Physical Connected Operator Solving Pressure Fields: Stage 2
This section will detail the second phase of our methodology: a PINN-Transformer architecture capable of reconstructing pressure fields from small sample data, featuring residual connections. In this architecture, the permeability field $\kappa$ and the multiscale basis functions $bf$ obtained from the first stage will serve as input data to learn the pressure solution $p(\kappa,bf)$ of Darcy equation. In other words, we aim to develop a parameterized deep learning model $\mathcal{G}$ that facilitates the mapping from input to output:
$\mathcal{G}:\ \mathbb{R}^{2\times 2}\mapsto\mathbb{R}^{2},\ \ \ (\kappa,bf)\mapsto p_{\theta}(\kappa,bf)$ (10)
@@ 4.1 Physics-informed Operator with Non-neighbored Accelerator
PGDL methods combine the powerful nonlinear mapping capabilities of deep learning with prior knowledge of physical laws, thereby significantly enhancing the accuracy, stability, and interpretability of predictive models. The PINN serves as a specific implementation of PGDL, embedding physical information into the model through residual learning.
@@
$\mathcal{R}(x;\theta)$ $=-\nabla\cdot\left(\kappa(x)\nabla p_{\theta}(x)\right)-f(x)$ (11)
$=-\left(\nabla\kappa(x)\cdot\nabla p_{\theta}(x)+\kappa(x)\Delta p_{\theta}(x)\right)-f(x)$ (12)
@@
To embed physical residuals into a neural network, we utilize a loss function to achieve this objective. Traditional neural networks typically employ the MSE loss for regression tasks or the Binary Cross-Entropy (BCE) loss for binary classification tasks. However, in our research, we need to incorporate physical information into the loss function to ensure that the model’s predictions adhere to the laws of physics. Additionally, we must consider the loss associated with the data itself to enhance the model’s fit to the actual observed data.
@@
$\mathcal{L}_{\text{data}}(\theta)=\frac{1}{N}\sum_{j=1}^{N}\|p_{\theta}(x_{j})-p_{\text{true}}(x_{j})\|^{2}$ (13)
@@
$\mathcal{L}_{\text{physics}}=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\|\mathcal{R}(x_{i};\theta)\|^{2}$ (15)
@@
Fig.4-stage2-local information extractor illustrates the operator architecture we propose, which consists of symmetrically distributed V-shaped convolutional layers and embeds physical information through residual learning during backpropagation. The increase in model complexity may lead to a higher risk of overfitting, which is undesirable. To address this, in addition to incorporating Dropout layers to mitigate the potential for overfitting, we have also introduced L2 regularization into the model, which adds a regularization term to the loss function:
@@
As the network architecture becomes increasingly complex, a significant loss of data features may occur during the transmission of information between layers, potentially leading to the gradient vanishing problem. Additionally, a large number of parameters can substantially prolong the training time of the model. To effectively expedite the training process, we have incorporated accelerators based on residual connections into the network architecture, aiming to mitigate these issues and enhance training efficiency.
@@
where $\mathcal{G}^{(b)}$ represents the original mapping function from layer (b1)(b-1) to layer $b$, a process that is independent of the skip connection.
These non-adjacent skip connections act as shortcuts from earlier layers to later layers in the computational graph, creating a multipath structure. During forward propagation, features from layer $a$ can directly influence the output of layer $b$ without undergoing the nonlinear transformations of the intermediate layers. This design ensures that the feature information extracted by the early layers is not easily diminished or lost during deep propagation.
Assume that the existence of a skip connection in the network that directly jumps from the input layer to the L-th layer: